You invested $100,000 into an account 25 years ago. The investment is now worth $350,000. Calculate the interest rate if the investment was compounded semi-annually.
Answer #1

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

350,000=100,000*(1+r/2)^(2*25)

(350,000/100,000)^(1/50)=(1+r/2)

(1+r/2)=1.02537178

r=(1.02537178-1)*2

=5.07%(Approx).